In Nyeri town there are only two milk processors. The local inverse demand for milk is given by: Q = 120− P, where P denotes price, Q denotes the total quantity measured in cartons. Both milk processors have the same cost function given by C = 30Q, where C is total cost and Q is output measured in cartons. What is the industry output (Q)?


1
Expert's answer
2022-01-18T09:59:22-0500

"Q=120-P"

Inverse Demand

"P=120-Q"

"C=30Q"

"MC=30"

"TR=P\times Q"

"=(120-Q)Q= 120Q-Q^2"

"MR=120-2Q"

For profit Maximization Out,

MR=MC

120-2Q=30

2Q=90

Q= 45